# Dense SAE Latents Are Features, Not Bugs

### v1

@@ 1 Introduction
Sparse autoencoders (SAEs) are an unsupervised method for extracting interpretable features from language models (bricken2023monosemanticity, ; huben2024sparse, ; kissane2024interpreting, ). They address the challenge of polysemanticity, where individual neurons activate in semantically diverse contexts that defy a single explanation (olah2017feature, ; elhage2022superposition, ). SAEs are trained to reconstruct the activations of a language model under a sparsity constraint applied to a bottleneck layer, ensuring that only a small subset of latents are active at a time. [^1] We use “latent” to refer to an entry in the SAE’s sparse hidden layer. This method effectively recovers interpretable features in a variety of models, including Claude 3 Sonnet (templeton2024scaling, ) and GPT-4 (gao2025scaling, ).
Ideally, a trained SAE would yield a large set of interpretable and sparsely activating latents. In practice, however, SAEs exhibit a substantial fraction of densely activating latents, activating on 10% to 50% of tokens (topkvsgated, ; rajamanoharan2024jumpingaheadimprovingreconstruction, ). These dense latents are challenging to interpret based solely on their activation patterns. It remains unclear whether they arise as an optimization by-product, or if they instead capture inherently dense signals present in the model’s residual stream (anthropic-interpretable-dense-features, ; removing-dense-latents, ).
In this work, we investigate several properties of dense SAE latents and the residual stream subspaces they span, uncovering evidence that these latents track meaningful residual stream information. First, we observe that when retraining an SAE on model activations with the dense latent space ablated, virtually no dense latents are learned—dense latents reflect an intrinsic property of the residual stream rather than a training artifact. We then study the geometry of dense latents and observe that they tend to form antipodal pairs, effectively reconstructing specific subspace directions.
We then examine the Gemma Scope suite of SAEs (lieberum2024gemmascopeopensparse, ) across layers to propose a taxonomy of dense latents. We identify latents whose activations encode positional information, latents reconstructing a subspace of the residual stream linked to entropy regulation (stolfo2024confidence, ; cancedda-2024-spectral, ), latents tracking high level shifts in the text, latents encoding letter-specific output signals, latents tracking parts of speech, and latents reconstructing the first residual stream principal component direction. We additionally examine how these dense latents transform across layers, finding that there is a pronounced increase in the number of dense latents just before the unembedding, as well as a shift from structural signals in early layers (e.g., position tracking) to output-oriented signals at the end. Our findings provide evidence that dense SAE latents reflect inherently dense mechanistic functions within language models.
@@ 2 Background
Sparse autoencoders (SAEs) are trained to reconstruct a language model’s activations $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{model}}}$ while imposing a sparsity constraint (yun-etal-2021-transformer, ; huben2024sparse, ). This computation can be represented as:
@@
We focus our investigation on the Gemma Scope SAEs (lieberum2024gemmascopeopensparse, ) trained on Gemma 2 2B (gemmateam2024gemma2improvingopen, ), which use a JumpReLU activation function (rajamanoharan2024jumpingaheadimprovingreconstruction, ). We additionally train TopK SAEs (gao2025scaling, ) on 1B tokens of the OpenWebText corpus (gokaslan2019openweb, ) for our experiments in § 3.1.222We choose TopK for its reliable training and competitive reconstruction–sparsity trade-off. Activation densities for Gemma Scope latents are from Neuronpedia (neuronpedia, ), while densities for our TopK SAEs are computed over 100M tokens from the C4 Corpus (c4, ). Full experimental details are in Appendix B.
@@ 3 General Properties of Dense Latents
We begin by examining structural properties of dense SAE latents, finding that they arise from a specific residual stream subspace (§ 3.1), and that they tend to cluster in antipodal pairs (§ 3.2).
@@ 3.1 Dense Latents Reflect Intrinsic Properties of the Residual Stream
To determine whether dense SAE latents arise from the training procedure or reflect an intrinsic property of the residual-stream subspace they reconstruct, we perform a targeted ablation experiment. We identify the subspace spanned by the dense latents of an SAE trained on layer 25 of Gemma 2 2B, then train a new SAE on activations in which this subspace has been zero-ablated. For comparison, we also select an equally sized set of non-dense latents and train a third SAE after ablating their subspace. We repeat this for two dictionary sizes ($d_{\mathrm{sae}}=16384$ and $32768$).
Figure 1a shows the resulting distributions of latent activation densities. In both dictionary sizes, ablating the dense-latent subspace (teal) yields much fewer high-density latents than the original SAE (blue) and the non-dense ablation (orange). This result implies that densely activating latents are not mere training artifacts but instead track a dense residual-stream subspace whose presence drives the emergence of dense latents. As additional evidence that dense latents are not training artifacts, in § A.2 we show that longer training does not reduce the number of dense latents.
@@ 3.2 Dense Latents Cluster in Antipodal Pairs
We now examine the geometry of dense latents and observe that they tend to form antipodal pairs. That is, as shown in Figure 1b, there exist many pairs of two dense latents that have nearly opposite decoder vectors (we find a similar result for encoder vectors). This suggests that the SAE allocates two latents in the dictionary to represent a 1-dimensional line.
@@
where $\operatorname{sim}(u,v)$ denotes the cosine similarity between vectors $u$ and $v$. This score reflects the extent to which latent $i$ forms an antipodal pairing with another latent: high values of $s_{i}$ indicate that there is another latent $j$ with both encoder and decoder weights nearly opposite in direction to those of $i$ [^3] Although high values of $s$ could be produced by two nearly identical latents, retaining such a pair would be redundant–a scenario we do not observe. Evidence for this is provided in § A.4.
As shown in Figure 1c, $s_{i}$ and the activation density of latent $i$ are strongly positively correlated. The majority of dense latents–particularly those with an activation frequency exceeding 0.3–exhibit pairwise scores greater than 0.9, supporting our conclusions above. We provide density-antipodality visualizations for additional SAEs in § A.1, showing that this trend holds consistently across SAE architectures (JumpReLU and TopK), models (GPT-2 and Gemma), and layers.
@@ 4 Taxonomy
  • Position latents (§ 4.1) fire based on token position relative to structural boundaries (start of sentence, paragraph or context) and appear early in the network.
  • Context-binding latents (§ 4.2) represent context-dependent semantic content and exhibit coherent chunk-level activations, potentially representing high-level ideas within the context.
  • Nullspace latents (Figure 5) track components of the residual stream that have minimal impact on next token prediction. They instead regulate prediction entropy.
  • Alphabet latents (§ 4.4) promote broad sets of tokens sharing an initial character.
  • Meaningful-word latents (§ 4.5) have activations related to the token part-of-speech tag.
  • PCA latents (§ 4.6) lie almost completely within the first PCA components of the activation space.
Figure 2: An overview of our taxonomy of dense latents, for every layer. See § E.1 for how we created this plot.
@@ 4.1 Position Latents
We first identify a class of dense latents whose activations track the current token’s position relative to specific text boundaries. Context-tracking latents track token position w.r.t. the BOS token, paragraph-tracking latents track token position w.r.t. a paragraph start, and sentence-tracking latents track token position w.r.t. a sentence beginning. Context-position latents are similar to “position neurons” from prior work (gurnee2024universalneuronsgpt2language, ); the other categories are to the best of our knowledge novel.
To find these latents systematically, we use Spearman’s rank correlation coefficent $\rho$. For each dense latent, we capture the projections [^4] We use the projection of the residual stream rather than the JumpRELU activations of these latents since we hypothesize that the direction itself encodes the positional information, regardless of whether the magnitude exceeds the learned JumpRELU threshold. of the residual stream activations onto its decoder vector for 5000 1024-token-long contexts. We find $\rho$ between this projection and the distance from the last period, the last newline, and the beginning of the input. These boundaries act as proxies for “beginning of sentence”, “beginning of paragraph” and “beginning of context”, respectively.
Figure 2 shows the resulting trends: sentence-tracking and paragraph-tracking latents are prominent before layer 10, while context-position tracking latents are present throughout the model. Figure 13 shows $\rho$ for all latents across layers. We can clearly see groups of outlier latents for each category, and thus classify latents as belonging to that category if $|\rho|>0.4$. Indeed, examples in § E.2 confirm that the identified outlier latents have position-tracking behavior. Notably, § E.2 also shows that paragraph-tracking latents are agnostic to artificially adding formatting newlines, suggesting that this direction in the model tracks true semantic paragraph breaks. Thus, our “distance to new line objective” is just a proxy. We also note that latents with high $\rho$ with periods also have high $\rho$ with newlines, since newlines and periods are correlated in text. In Figure 15, we thus show the $\rho$ for sentence-tracking vs. paragraph-tracking across all dense latents.
@@ 4.2 Context-Binding Latents
Table 1: Fraction of “unclear” judgements using in-context examples versus out-of-context examples, for the highest-scoring latents by flips.
We next identify a class of dense latents that encode different semantic concepts depending on context. Unlike interpretable sparse SAE latents typically associated with fixed meanings, such as the “Golden Gate Bridge” feature in Claude (templeton2024scaling, ), these dense latents appear to bind to the main ideas of the context.
We first observe that some dense latents, particularly in middle layers, activate on long consecutive “chunks” of tokens. [^5] While positional latents also exhibit consecutive activations, here we refer to non-positional latents whose activations cannot be explained by position alone. Examining the activations of such latents and attempting to explain them with an LLM (see § E.5), we notice empirically that such latents fire on highly specific concepts within a context, but the concepts vary across contexts. One possible interpretation is that these latents represent general but abstract, difficult-to-interpret properties. However, we also observe that within an antipodal pair, the active latent often switches when the main topic or entity in the text changes (Figure 3, § E.3). This raises the hypothesis that such directions act as a “registers” in the residual stream for tracking the active concept, rather than simply representing generic properties.
We thus perform a steering experiment to find the causal effect of these directions. For each antipodal pair (F1, F2), we prompt Gemma 2 2B with input text from the RedPajama dataset (weber2024redpajama, ) and generate completions without steering, steering to F1, and steering to F2. An LLM judge (google2025gemini, ) is then asked whether each completion is more in line with activating examples (from the input context) of F1 or F2, or unclear. Further details of the methodology are in § E.4.
Since the unsteered generation may already favor F1 or F2, we quantify steering success by the fraction of flips from the unsteered judgement that align correctly with the steering direction. For several mid-layer latent pairs, steering reliably shifts completions toward the specific concept previously associated with the latent in that context. However, when judged against out-of-context examples, the rate of unclear judgements rises sharply. While difficult to rule out the possibility that these directions encode “general uninterpretable” features, the specificity of the steered generation in bringing up context-related ideas (Figure 3) suggests that these latents could bind to concepts in a context-dependent, rather than globally consistent, way.
Previous works have uncovered “binding mechanisms” that help the model keep track of in-context associations between entities (jiahai_binding, ; feng2024monitoringlatentworldstates, ). While our findings do not directly prove such a mechanism, they raise the possibility that dense subspaces may play a similar functional role, distinguishing the currently active semantic concept. Further work could explore the circuits (marks2025sparse, ) involving such subspaces, and challenge the assumption of globally monosemantic directions.
@@ 4.3 Nullspace Latents
Previous work has identified a $\mathbf{W}_{\mathrm{U}}$ quasi-nullspace–the subspace spanned by the last singular vectors of the unembedding matrix $\mathbf{W}_{\mathrm{U}}$–which accounts for a substantial portion of the residual stream’s norm, yet has little direct impact on next-token prediction (cancedda-2024-spectral, ). Since this subspace carries high norm, we hypothesize that some dense SAE latents are allocated specifically to reconstruct it.
@@
A histogram of $\alpha_{10}$ for the SAE trained at layer 25 of Gemma 2 2B (Figure 5a) shows that $99.6\%$ of latents have $\alpha_{10}<0.2$. We designate those with $\alpha_{10}>0.2$ as nullspace-aligned. Interestingly, $75\%$ of them are high-density, and account for $40\%$ of the high-density latents in the SAE.
Unlike other dense latents, nullspace-aligned latents are hard to interpret via their token-level activation patterns. Additionally, the tokens they promote are typically uninterpretable “under-trained” tokens (land-bartolo-2024-fishing, ). Motivated by prior work linking the $\mathbf{W}_{\mathrm{U}}$ nullspace to an RMSNorm-based (rmsnorm, ) entropy regulation mechanism (stolfo2024confidence, ), we investigate whether these latents encode this internal computation.
To test whether these latents causally influence output entropy, we ablate the residual stream along each latent’s decoder direction by setting its value to the corresponding decoder bias, thereby removing information in that direction. We then measure the change in per-token entropy of the model’s output distribution. Figure 5b reports the entropy change for all latents with $\alpha_{10}>0.3$ (one per antipodal pair to avoid redundancy), compared to a control group of 50 randomly selected latents. [^6] The entropy changes for the random latents are aggregated into a single boxplot.
We find that some nullspace latents produce much larger entropy shifts than the random baseline, indicating that they encode signals relevant to entropy modulation. In particular, latent 14325 has a disproportionate impact on output entropy. To test whether this signal is used by the model in conjunction with RMSNorm scaling (as in (stolfo2024confidence, )), we repeat the ablation while freezing the RMSNorm scaling coefficient. Figure 5c shows that the entropy change diminishes under this intervention, suggesting that the model uses this direction to modulate entropy via RMSNorm. Furthermore, Figure 6 shows that the combined activation of the antipodal pair formed by latents 13748 and 14325 is strongly correlated with output entropy, further supporting this interpretation.
While these results highlight the functional role of specific nullspace latents in entropy regulation, not all latents in this subspace behave similarly. Some exhibit negligible impact on entropy when ablated. We speculate that these may track different internal signals–one such candidate is the attention sink signal, which has also been associated with the $\mathbf{W}_{\mathrm{U}}$ nullspace (cancedda-2024-spectral, ). Overall, these experiments provide mechanistic evidence that nullspace latents correspond to internal model computations.

### v2

@@ 1 Introduction
Sparse autoencoders (SAEs) are an unsupervised method for extracting interpretable features from language models (bricken2023monosemanticity; huben2024sparse; kissane2024interpreting). They address the challenge of polysemanticity, where individual neurons activate in semantically diverse contexts that defy a single explanation (olah2017feature; elhage2022superposition). SAEs are trained to reconstruct the activations of a language model under a sparsity constraint applied to a bottleneck layer, ensuring that only a small subset of latents are active at a time. [^1] We use “latent” to refer to an entry in the SAE’s sparse hidden layer. This method effectively recovers interpretable features in a variety of models, including Claude 3 Sonnet (templeton2024scaling) and GPT-4 (gao2025scaling).
Ideally, a trained SAE would yield a large set of interpretable and sparsely activating latents. In practice, however, SAEs exhibit a substantial fraction of densely activating latents, activating on 10% to 50% of tokens (topkvsgated; rajamanoharan2024jumpingaheadimprovingreconstruction). These dense latents are challenging to interpret based solely on their activation patterns. It remains unclear whether they arise as an optimization by-product, or if they instead capture inherently dense signals present in the model’s residual stream (anthropic-interpretable-dense-features; removing-dense-latents).
In this work, we investigate several properties of dense SAE latents and the residual stream subspaces they span, uncovering evidence that these latents track meaningful residual stream information. First, we observe that when retraining an SAE on model activations with the dense latent space ablated, virtually no dense latents are learned—dense latents reflect an intrinsic property of the residual stream rather than a training artifact. We then study the geometry of dense latents and observe that they tend to form antipodal pairs, with each pair effectively reconstructing a single direction.
We then examine the Gemma Scope suite of SAEs (lieberum2024gemmascopeopensparse) across layers to propose a taxonomy of dense latents. We identify latents whose activations encode positional information, latents reconstructing a subspace of the residual stream linked to entropy regulation (stolfo2024confidence; cancedda-2024-spectral), latents tracking high-level shifts in the text, latents encoding letter-specific output signals, latents tracking parts of speech, and latents reconstructing the first residual stream principal component direction. We additionally examine how these dense latents transform across layers, finding that there is a pronounced increase in the number of dense latents just before the unembedding, as well as a shift from structural signals in early layers (e.g., position tracking) to output‐oriented signals at the end. Our findings provide evidence that dense SAE latents reflect inherently dense mechanistic functions within language models.
@@ 2 Background
Sparse autoencoders (SAEs) are trained to reconstruct a language model’s activations $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{model}}}$ while imposing a sparsity constraint (yun-etal-2021-transformer; huben2024sparse). This computation can be represented as:
@@
We focus our investigation on the Gemma Scope SAEs (lieberum2024gemmascopeopensparse) trained on Gemma 2 2B (gemmateam2024gemma2improvingopen), which use a JumpReLU activation function (rajamanoharan2024jumpingaheadimprovingreconstruction). We additionally train TopK SAEs (gao2025scaling) on 1B tokens of the OpenWebText corpus (gokaslan2019openweb) for our experiments in §˜3.1.222We choose TopK for its reliable training and competitive reconstruction–sparsity trade-off. Activation densities for Gemma Scope latents are from Neuronpedia (neuronpedia), while densities for our TopK SAEs are computed over 100M tokens from the C4 Corpus (c4). Full experimental details are in Appendix˜B.
@@ 3 General Properties of Dense Latents
We begin by examining structural properties of dense SAE latents, finding that they arise from a specific residual stream subspace (§˜3.1), and that they tend to cluster in antipodal pairs (§˜3.2).
@@ 3.1 Dense Latents Reflect Intrinsic Properties of the Residual Stream
To determine whether dense SAE latents arise from the training procedure or reflect an intrinsic property of the residual‐stream subspace they reconstruct, we perform a targeted ablation experiment. We identify the subspace spanned by the dense latents of an SAE trained on layer 25 of Gemma 2 2B, then train a new SAE on activations in which this subspace has been zero-ablated. For comparison, we also select an equally sized set of non‐dense latents and train a third SAE after ablating their subspace. We repeat this for two dictionary sizes ($d_{\mathrm{sae}}=16384$ and $32768$).
Figure˜1a shows the resulting distributions of latent activation densities. In both dictionary sizes, ablating the dense‐latent subspace (teal) yields much fewer high‐density latents than the original SAE (blue) and the non‐dense ablation (orange). This result implies that densely activating latents are not mere training artifacts but instead track a dense residual‐stream subspace whose presence drives the emergence of dense latents. As additional evidence that dense latents are not training artifacts, in §˜A.2 we show that longer training does not reduce the number of dense latents. We further replicate this dense‐subspace ablation on GPT-2 and LLaMA 3.2 with the same outcome (§˜A.6).
@@ 3.2 Dense Latents Cluster in Antipodal Pairs
We now examine the geometry of dense latents and observe that they tend to form antipodal pairs. That is, as shown in Figure˜1b, there exist many pairs of two dense latents that have nearly opposite decoder vectors (we find a similar result for encoder vectors). This suggests that the SAE allocates two latents in the dictionary to represent a 1-dimensional line.
@@
Figure 2: AbsoluteTopK SAEs show no antipodality. Allowing the SAE to have both positive and negative latent activations removes antipodal dense latents.
where $\operatorname{sim}(u,v)$ denotes the cosine similarity between vectors $u$ and $v$. This score reflects the extent to which latent $i$ forms an antipodal pairing with another latent: high values of $s_{i}$ indicate that there is another latent $j$ with both encoder and decoder weights nearly opposite in direction to those of ii. [^3] Although high values of $s$ could be produced by two nearly identical latents, retaining such a pair would be redundant–a scenario we do not observe. Evidence for this is provided in § A.4.
As shown in Figure˜1c, $s_{i}$ and the activation density of latent $i$ are strongly positively correlated. The majority of dense latents—particularly those with an activation frequency exceeding 0.3—exhibit pairwise scores greater than 0.9, supporting our conclusions above. We provide density-antipodality visualizations for additional SAEs in §˜A.1, showing that this trend holds consistently across SAE architectures (JumpReLU and TopK), models (GPT-2 and Gemma), and layers.
Additionally, we train an AbsoluteTopK SAE, which allows activations of SAE latents to be negative, and enforces sparsity by taking the TopK latents with greatest absolute activations. This effectively allows the same latent direction to be used in both “positive” and “negative” directions for reconstruction. We compare this to a TopK SAE trained with the same seed, and show that this AbsoluteTopK activation function eliminates the antipodal dense latents (Figure 2).
@@ 4 Taxonomy
  • Position latents (§˜4.1) fire based on token position relative to structural boundaries (start of sentence, paragraph or context) and appear early in the network.
  • Context-binding latents (§˜4.2) represent context-dependent semantic content and exhibit coherent chunk-level activations, potentially representing high-level ideas within the context.
  • Nullspace latents (§˜4.3) track components of the residual stream that have minimal impact on next token prediction. They instead regulate prediction entropy.
  • Alphabet latents (§˜4.4) promote broad sets of tokens sharing an initial character.
  • Meaningful-word latents (§˜4.5) have activations related to the token part-of-speech tag.
  • PCA latents (§˜4.6) lie almost completely within the first PCA components of the activation space.
Figure 3: An overview of our taxonomy of dense latents, for every layer. See §˜E.1 for how we created this plot.
@@ 4.1 Position Latents
We first identify a class of dense latents whose activations track the current token’s position relative to specific text boundaries. Context-tracking latents track token position w.r.t. the BOS token, paragraph-tracking latents track token position w.r.t. a paragraph start, and sentence-tracking latents track token position w.r.t. a sentence beginning. Context-position latents are similar to “position neurons” from prior work (gurnee2024universalneuronsgpt2language); the other categories are to the best of our knowledge novel.
To find these latents systematically, we use Spearman’s rank correlation coefficient $\rho$. For each dense latent, we capture the projections [^4] We use the projection of the residual stream rather than the JumpReLU activations of these latents since we hypothesize that the direction itself encodes the positional information, regardless of whether the magnitude exceeds the learned JumpReLU threshold. of the residual stream activations onto its decoder vector for 5000 1024-token-long contexts. We find $\rho$ between this projection and the distance from the last period, the last newline, and the beginning of the input. These boundaries act as proxies for “beginning of sentence”, “beginning of paragraph”, and “beginning of context”, respectively.
Figure˜3 shows the resulting trends: sentence-tracking and paragraph-tracking latents are prominent before layer 10, while context-position tracking latents are present throughout the model. Figure˜15 shows $\rho$ for all latents across layers. We can clearly see groups of outlier latents for each category, and thus classify latents as belonging to that category if $|\rho|>0.4$. Indeed, examples in §˜E.2 confirm that the identified outlier latents have position-tracking behavior. Notably, §˜E.2 also shows that paragraph-tracking latents are agnostic to artificially adding formatting newlines, suggesting that this direction in the model tracks true semantic paragraph breaks. Thus, our “distance to newline objective” is just a proxy. We also note that latents with high $\rho$ with periods also have high $\rho$ with newlines, since newlines and periods are correlated in text. In Figure˜17, we thus show the $\rho$ for sentence-tracking vs. paragraph-tracking across all dense latents.
@@ 4.2 Context-Binding Latents
We next identify a class of dense latents that encode different semantic concepts depending on context. Unlike interpretable sparse SAE latents typically associated with fixed meanings, such as the “Golden Gate Bridge” feature in Claude (templeton2024scaling), these dense latents appear to bind to the main ideas of the context.
We first observe that some dense latents, particularly in middle layers, activate on long consecutive “chunks” of tokens. [^5] While positional latents also exhibit consecutive activations, here we refer to non-positional latents whose activations cannot be explained by position alone. Examining the activations of such latents, we notice empirically that such latents fire on highly specific concepts within a context, but the concepts vary across contexts. We generate explanations of these latents with an LLM and confirm that they seem to be more context-specific than sparse latents (see §˜E.3).
One possible interpretation is that these latents represent general but abstract, difficult-to-interpret properties. However, we also observe that within an antipodal pair, the active latent often switches when the main topic or entity in the text changes (Figure˜4, §˜E.4). This raises the hypothesis that such directions act as “registers” in the residual stream for tracking the active concept, rather than simply representing generic properties.
We thus perform a steering experiment to find the causal effect of these directions. For each antipodal pair (F1, F2), we prompt Gemma 2 2B with input text from the RedPajama dataset (weber2024redpajama) and generate completions without steering, steering to F1, and steering to F2. An LLM judge (google2025gemini) is then asked whether each completion is more in line with activating examples (from the input context) of F1 or F2, or unclear. Further details of the methodology are in §˜E.5.
Since the unsteered generation may already favor F1 or F2, we quantify steering success by the fraction of flips from the unsteered judgment that align correctly with the steering direction. For several mid-layer latent pairs, steering reliably shifts completions towards the specific concept previously associated with the latent in that context (Figure˜5). However, when judged against out-of-context examples, the rate of unclear judgments rises sharply (Table˜1). While difficult to rule out the possibility that these directions encode “general uninterpretable” features, the specificity of the steered generation in bringing up context-related ideas suggests that these latents could bind to concepts in a context-dependent, rather than globally consistent, way.
Previous works have uncovered “binding mechanisms” that help the model keep track of in-context associations between entities (jiahai_binding; feng2024monitoringlatentworldstates). While our findings do not directly prove such a mechanism, they raise the possibility that dense subspaces may play a similar functional role, distinguishing the currently active semantic concept. Further work could explore the circuits (marks2025sparse) involving such subspaces, and challenge the assumption of globally monosemantic directions.
@@
Table 1: Fraction of “unclear” judgments using in-context examples versus out-of-context examples, for the highest-scoring latents by flips.
@@ 4.3 Nullspace Latents
Previous work has identified a $\mathbf{W}_{\mathrm{U}}$ quasi-nullspace–the subspace spanned by the last singular vectors of the unembedding matrix $\mathbf{W}_{\mathrm{U}}$–which accounts for a substantial portion of the residual stream’s norm, yet has little direct impact on next‐token prediction (cancedda-2024-spectral). Since this subspace carries high norm, we hypothesize that some dense SAE latents are allocated specifically to reconstruct it.
@@
A histogram of $\alpha_{10}$ for the SAE trained at layer 25 of Gemma 2 2B (Figure˜6a) shows that $99.6\%$ of latents have $\alpha_{10}<0.2$. We designate those with $\alpha_{10}>0.2$ as nullspace‐aligned. Interestingly, $75\%$ of them are high-density, and account for $40\%$ of the high-density latents in the SAE.
Unlike other dense latents, nullspace‐aligned latents are hard to interpret via their token‐level activation patterns. Additionally, the tokens they promote are typically uninterpretable “under‐trained” tokens (land-bartolo-2024-fishing). Motivated by prior work linking the $\mathbf{W}_{\mathrm{U}}$ nullspace to an RMSNorm-based (rmsnorm) entropy regulation mechanism (stolfo2024confidence), we investigate whether these latents encode this internal computation.
To test whether these latents causally influence output entropy, we ablate the residual stream along each latent’s decoder direction by setting its value to the corresponding decoder bias, thereby removing information in that direction. We then measure the change in per-token entropy of the model’s output distribution. Figure˜6b reports the entropy change for all latents with $\alpha_{10}>0.3$ (one per antipodal pair to avoid redundancy), compared to a control group of 50 randomly selected latents. [^6] The entropy changes for the random latents are aggregated into a single boxplot.
We find that some nullspace latents produce much larger entropy shifts than the random baseline, indicating that they encode signals relevant to entropy modulation. In particular, latent 14325 has a disproportionate impact on output entropy. To test whether this signal is used by the model in conjunction with RMSNorm scaling (as in stolfo2024confidence), we repeat the ablation while freezing the RMSNorm scaling coefficient. Figure˜6c shows that the entropy change diminishes under this intervention, suggesting that the model uses this direction to modulate entropy via RMSNorm. Furthermore, Figure˜7 shows that the combined activation of the antipodal pair formed by latents 13748 and 14325 is strongly correlated with output entropy, further supporting this interpretation.
While these results highlight the functional role of specific nullspace latents in entropy regulation, not all latents in this subspace behave similarly. Some exhibit negligible impact on entropy when ablated. We speculate that these may track different internal signals–one such candidate is the attention sink signal, which has also been associated with the $\mathbf{W}_{\mathrm{U}}$ nullspace (cancedda-2024-spectral). Overall, these experiments provide mechanistic evidence that nullspace latents correspond to internal model computations.